Helen’s preferences over CDs (C) and sandwiches (S) are given by U(S,C) = SC + 10(S + C),with MUC = S+ 10 and MUS =C+ 10.If the price of a CD is $9 and the price of a sandwich is $3,and Helen can spend a combined total of $30 each day on these goods, find Helen’s optimal consumption basket.
1
Expert's answer
2020-12-04T08:02:55-0500

First of all we will say

Budget"30=9C+3S or S=10-3C"

Therefore

"U = S\u2022C+10(S+C) = S\u2022C+10S+10C = C(10-3C)+10(10-3C)+10C"

Simplify to

"U = 10C-3C\u00b2+100-30C+10C = 100-10C-3C\u00b2"

this means

"C\u22650 and from budget: C[MAX]=30\/9\u22483.33 ; at S=0"


"S\u22650 and from budget: S[MAX]=30\/3=10.00 ; at C=0"


"3.33\u2265C\u22650"


"0\u2264S\u226410"


Now maximize U(C)→MAX

100-10C-3C²→MAX & 10/3≥C≥0

∂U(C)/∂C= -10-3C < 0

Maximum will be a lower boundary since its downward sloping for all allowed ranges.

Therefore utility-maximizing solution backet is:C=0 & S=10 &U[MAX]=100.